A Complete Bibliography of Publications in Ecological Modelling (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.02

Title word cross-reference

$14.50 [DM75]. $16.00 [Smi77]. $16.50 [Bis78]. $17.50 [Jør76c]. $18.75 [Jør76a]. $19.00 [Nih78]. $19.50 [Jør75b]. $20 [Ren79]. $20.00 [Jør77b]. $22.45 [Mej79]. $23.00 [Rey77]. $24.75 [Gro79]. $24.80 [Sko79]. $25.85 [Jac79]. $27.75 [Jør77d]. $30.75 [Jør79c]. $33.95 [Mid79]. $35 [Ren79]. $38.00 [Jør79b]. $47.95 [Jør77e]. $49.95 [Mor79]. ^{32} [HUS79]. ^{60} [BH77]. x [Mej79].

0 [Bis78, Jac79, Jes76, Jør76a, Jør76c, Jør77d, Jør77e, Jør79c, Jør79a, Mid79, Mor79, Nih78, Nih79, Smi77]. 0-08-020611-5 [Nih78]. 0-12-547204-8 [Jør79a]. 0-250-40077-4 [Jac79]. 0-275-24190-4 [Bis78]. 0-444-00157-3 [Mid79]. 0-444-41259-X [Jør76a]. 0-444-41270-0 [Jør77e]. 0-444-41313-8 [Jør76c]. 0-444-41329-4 [Jør79c]. 0-444-41419-3 [Jes76]. 0-444-41560-2 [Mor79]. 0-470-94361-0 [Jør77d]. 0-471-68925-4 [Smi77]. 0-471-99580-0 [Nih79]. 0471-01993 [Mej79].

17th [BB75]. 1975 [Nih78].

2 [Mor79]. 20.00 [Jac75]. 250 [Gro79]. 250-401045 [Gro79].

3 [Jes76, Jør77a, Mej79, Mid79, Sko79]. 3-540-90269-4 [Sko79]. 3.70 [Jør75d].

4 [Bis78, Jac79, Jør79c, Sko79, Smi77]. 401045 [Gro79].

5 [Nih78].

6.40 [Els76].

70.00 [Jør75a].

8 [Jør76c, Jør79a]. 8th [BB75].

9.80 [Con75]. 90 [Jør77a]. 90-220-0593-3 [Jør77a].

abundance [SLM^{+}79]. Academic [Jør79a]. acarine [Jac75]. accommodation [BS78]. Adequacy [HR78]. Advanced [Nih78]. advection [HL78]. aerobic [KN75]. affected [McK77]. Age [Gop77]. Aggregation [OR79]. Agricultural [Jør77a]. Alan [Con75]. algae [NW79, SLM^{+}79, Sei79]. algal [DLvS78, FKBB75]. algorithms [BH77]. Alice [Mit76]. allocation [HBJ76]. Amsterdam [Jes76, Jør75a, Jør76a, Jør76c, Jør77e, Jør79c, Mid79, Mor79]. Amsterdam-Oxford-New [Mor79]. anaerobic [JKNJ75]. analyses [HR79]. Analysis [Mag75, Bar78a, Bar79, BH77, Cri78, DW78, EWS78, Jør77a, Jør79a, Les76, Ren79, Foe76]. Analytical [Hal77]. animal [Boy79]. Ann [Con75, DM75, Gro79, Jac79]. Announcement [Ano76a, Ano77a, Ano77b, Ano78a, Ano78b, Ano79a, Ano79b, Ano79c, Ano79d]. Announcements [Ano78c, Ano79e]. Application [CBB77, DLB78b, GR77, HUS79]. Applied [Foe76, Nih76]. Applying [Kin78]. approach [BB75, JM79, KN75, OFA75, PFF76, Rey77, Jør79c]. approaches [Jør79g]. aquatic [BH77, BBMT75, Gro79, Jør79e, SP76, Str79, THR^{+}78]. Aqueous [Con75]. Aqueous-environmental [Con75]. Arbor [Con75, DM75, Gro79, Jac79]. arid [Kri78]. Arie [Jes76]. Arnold [Jør77a]. art [Ano76i]. Aspects [Arn78, CHPS78, BBMT75, Jør75b]. assembled [GJ77]. assessing [WM76]. assessment [GR77]. associated [Hal77]. Association [Ano77h]. atmospheric [GWJ79, WGWJ79]. attention [Gil77]. August [Ano77k]. Author [Ano75a, Ano76b, Ano77c, Ano78d, Ano78e, Ano79f, Ano79g].

B [Jør79c, Nih78]. balance [Cri78, HM78, JM76a, JM76b]. barren [WHP75]. barren-ground [WHP75]. Barrie [Jør79c]. Bay [OFA75, YU79, McK77]. behavior [DL77]. behaviour [BBMT75]. belowground [Gil77]. Benchmark [Jør77d]. benthic [SLM^{+}79, Sei79]. Berlin [Sko79]. Bernard [Jør79a]. between [Kam78]. Bight [Nih76]. bio [BS78]. bio-species [BS78]. biochemical [Gro79]. biodegradation [BV78a]. biological [Els76, VLG77]. biomass [DLB78a, DLB78b]. biostimulation [FKBB75]. biphenyls [KK79]. birth [Fra79]. Biswas [Jør77b, Rey77]. Biwa [IA78]. bluegill [CBB77]. Board [Ano75b, Ano76c, Ano77e, Ano78f, Ano78g, Ano79h, Ano79i]. BOD [PP79]. boll [JBB^{+}75]. books [Ano75h]. both [Ren79]. boundary [OFA75]. Britain [DM75]. British [Clo78]. buffer [JM77]. Bugliarello [Jør76a].

C [Jør75a, Jør77a, Jør79a, Mej79, Smi77]. Calculation [VS79]. California [Ren79]. call [Jør76d]. Cambridge [Jør75d]. Canale [Gro79]. capacity [JM77]. carbon [BV78b, JM76a]. caribou [WHP75]. case [CJH79, Jør79b, SW75, Zuc75]. Centre [Jør77a]. CERES [DLB78a, DLB78b]. chance [HBJ76]. chance-constrained [HBJ76]. Chapman [Mor79]. Charles [Jør79b]. chemical [Kin78]. chemistry [Con75]. Chichester [DM75]. circulations [Led79]. city [DL77]. cleaner [SP76]. closed [Par78a]. Co [BH77]. coastal [McK77, Mor79]. coefficients [SE76, VS79]. Coleoptera [JBB^{+}75]. colony [RM79]. COLTS [Jør76d]. COLTS-continuous [Jør76d]. Columbia [Clo78]. combining [Zuc75]. comments [HJ76]. Committee [Bis78]. communities [MP77]. Company [Rey77, Jes76]. compartmental [BH77]. competition [AAV78, MP77]. complex [Goh75]. Computer [WHP75, HJ76, Kal78, Jør76a, Mid79]. Computing [Ano77h]. concentration [WVW78]. concerning [Bar79]. Conference [Ano77d]. coniferous [Get78]. constituents [PFF76]. constrained [HBJ76]. constructs [SP76]. contaminant [DLB78a, DLB78b]. contamination [Jør76e]. continuous [Jør76d]. Control [VLG77, BD78, Jac79, Str79]. Cooperation [Jør79d]. coral [MP77]. Councils [Ren79]. Critical [Jør77a]. crops [Jør76e]. crude [SW75]. Cryptomonas [Clo78]. Curculionidae [JBB^{+}75]. cycle [IA76, JM76a]. cycling [Nyh78].

D [DM75]. data [CBB77]. David [Nih79]. Day [Jør79b]. December [Ano75f, Ano76g]. decision [Bis75a, Bis78]. decision-making [Bis75a]. decisions [PLM75]. decomposition [DW78]. Dekker [Jør75b]. density [Jac75]. dependent [Gop77]. depletion [VA76]. description [DLB78a]. design [CJH79]. Developments [Jør76a, Jør76c, Jør77e]. Dfl [Jac75, Jes76, Jør75a, Jør76a, Jør76c, Jør77a, Jør77e, Jør79c, Mid79, Mor79]. diagrams [GR77]. differential [HS77, THR^{+}78]. diffusion [HL78, Hil79]. directions [Ren79]. dispersed [Fra79]. dispersion [AA78, DLvS78, Gop77]. dissolved [WVW78]. distributed [DM75]. distribution [Jør79f, Jør79e, KK79, SLM^{+}79]. diurnal [Kam78]. diversity [KM79, Smi77]. DM [Sko79]. DO [PP79]. Documentation [Jør77a, SP76, Swa79]. Dowden [Jør77d]. Dubrovnik [BB75]. Dynamic [VA76, YU79, Ben79, IA78, IG78]. Dynamics [IA76, CKDF79, Clo78, DLB78a, DLB78b, Hil79, Kal78, Kri78, LN78, Mag75, NW79, Phi79, WHP75].

Ecological [Foe76, JM77, Sko79, Ano77d, CJH79, Jør75c, Jør76e, Jør78a, Jør78b, Jør79d, JM79, Jør79g, Kin78, Mag75, OR79, Ren79, Jør76b, Jør77d, Smi77]. ecology [Mej79, Jør77d, Jør75d, Jør79a, Mej79]. economic [SW75, Zuc75]. Ecosystem [Mit76, Bar78a, Bar78b, Bar79, CJH79, Gil77, Hal76, HR78, HR79, Jør79e, Mag75, McK77, MKU79, Par75, Par78a, Swa79, THR^{+}78, UFHH78, Jør79b]. Ecosystems [Mor79, BBMT75, CKDF79, Get78, Goh75, HS77, Jør77a, Kal78, KK79, Per78, Str79, Gro79, Mor79]. edition [Mej79]. Editor [Con75, Gro79, Jør75a, Jør75b, Jør77d, Jør77b, Jør79a, Mor79, Nih78, Ren79, Rey77, Sko79]. editorial [Jør76b, Ano75b, Ano76c, Ano77e, Ano78f, Ano78g, Ano79h, Ano79i]. Editors [DM75, Jac79, Jør77a, Jør79b, Nih79, Jes76]. effect [FKBB75, Jør79f, Jør79e]. effects [DLB78a, DLB78b, DW78, HL78, KK79]. Efficient [BH77]. Eigengroup [HR79]. Elaboration [DW78]. Elsevier [Jes76, Jør75a, Jør76a, Jør76c, Jør77e, Jør79c, Mid79, Mor79]. empirical [UFHH78]. employing [Agn79]. energetics [Jør77d]. Energy [Zuc75, Cri78, Swa79]. Energy-economic [Zuc75]. entropy [Mau79]. Environment [Bis78, Jør79c, Els76]. Environmental [Mid79, Par75, Bis75a, BB75, Foe76, GR77, Nih79, PLM75, Pet77, Bis78, Con75]. equation [Hal77, HS77, THR^{+}78]. equations [GWJ79]. Errata [Ano79j]. Erratum [Ano76d, Ano77f, Ano77g, Ano78h]. error [OR79]. Esrom [KN75]. Estimation [Fra79, BH77, HUS79, LN78]. estuaries [AA78]. estuarine [McK77, OFA75, PFF76]. estuary [Kal78]. ethological [Jør79c]. eutrophic [Nyh78]. eutrophication [DM75, HL78, IA78, Jør76f]. evaluation [Jør77a, Pet77]. event [Ano75c]. Everglades [BO76]. evolving [Per78]. Examination [JMF78]. example [HL78]. Examples [Pet77]. exchange [JKNJ75, Jør78b, KN75]. exhaustible [SW75]. exploitation [Agn79]. exponential [DG79].

F [Jør76a]. Falkenborg [DM75]. Feasibility [GJ77]. February [Ano77j, Ano78n, Ano79p]. feeding [JBB^{+}75, VS79]. field [Jør78b, Jør79d]. figs [Jac79]. figures [Jør75d, Jør77a, Jør79c, Mor79, Nih78, Nih79]. fire [BO76]. fires [Day77]. fish [DG79, Jør76g]. fishery [Agn79, Goh77]. fishing [Goh77]. fitting [Ben79]. fjord [SLM^{+}79]. Fleming [Mid79]. Florida [Mit76, Zuc75]. flow [Bar78a, Bar78b, Bar79, Swa79]. Focus [Jør78a]. forage [HBJ76]. forest [Day77, DLB78a, DLB78b, EWS78, Get78, MKU79, Pet77, Phi79]. forest-insect [Pet77]. Formation [Ano77h]. formulations [SE76]. Forthcoming [Ano75c]. Francisco [Jør79a]. Francois [Nih79]. Franz [Jac75]. Freedman [Jør76c]. Frenkiel [Nih79]. fresh [WVW78]. function [Agn79]. functional [Jac75].

G [Jac75, Jør76a, Jør77a, Jør77d, Ren79]. general [KM79]. George [Mid79, Sko79]. global [JM76a, JM76b]. Golterman [Jør77e]. Goodall [Nih79]. Graphical [Pet77]. grassland [Cal79, Gil77, Sko79]. grazing [Kri78, SE76, VS79]. Great [DM75]. Greenbie [Jør79c]. ground [Jør76e, WHP75]. growth [DLvS78, FKBB75, HJ76, Jør75e, Jør76g, MP77, Nyh78]. guide [Jes76]. Gunther [Jør76a].

H [DM75, Jac75, Jør76c, Jør77e]. habitat [WM76]. habitats [Gop77]. Haimes [Jør76c]. Hall [Jør76c, Jør79b]. Harvest [Jør77b]. harvesting [Agn79]. heat [JM76a, JM76b]. heavy [GWJ79, Jør76e, Jør79e, Sei79, WGWJ79]. Heidelberg [Sko79]. Herbage [Kri78]. herds [Boy79]. Hill [Rey77]. histories [Jør79b]. Holcomb [Bis78]. holistic [Gil77, JM79]. House [Jør77b]. Huntington [YU79]. Hutchinson [Jør77d]. hydrodynamical [HB78]. hydrological [CHPS78]. hydrology [Mid79]. hypothetical [HL78].

identification [HUS79]. II [DLB78b, GWJ79, Jør77b]. illustrations [Jør76a, Jør76c, Jør77d, Jør77b, Jør77e]. impact [GR77]. implications [SW75]. Improvement [Jør78b]. incorporation [LBD78]. index [Ano75a, Ano76b, Ano77c, Ano78d, Ano78e, Ano79f, Ano79g]. Indian [Kri78]. Indirect [DW78]. induced [DLvS78]. influence [JM76a]. inherently [Sjö77]. Innis [Ren79, Sko79]. Input [BD78]. insect [Pet77]. Institute [Bis78, Foe76, Nih78]. interaction [Kam78]. Interdisciplinary [Sha76, Nih76]. internal [Kam78]. international [Les76, Ano77h, Foe76]. interrelations [BO76]. Interscience [Els76]. intraspecific [AAV78]. intrinsic [HS77]. introduction [Jør79b, Mej79, Per78]. invertebrates [DW78]. Irish [HB78]. ISBN [Bis78, Gro79, Jac79, Jes76, Jør76a, Jør76c, Jør77a, Jør77d, Jør77e, Jør79c, Jør79a, Mej79, Mid79, Mor79, Nih78, Nih79, Sko79, Smi77]. issues [PLM75]. Italy [Nih78]. IV [Jør79a].

J [Con75, DM75, Els76, Jør75d, Jør75a, Mor79]. Jamaica [OFA75]. January [Ano78i, Ano79k]. John [Jør75b, Jør79b, Nih79, Smi77]. Jolla [Ren79]. journal [Jør76b, Jør77c]. Jr [Jør79b]. July [Ano75g, Ano78o, Ano79q]. June [Ano76h, Ano78j, Ano79l]. juvenile [Fra79]. juveniles [Fra79].

Keinath [Jac79]. kenetics [HUS79]. kinematic [DL77]. kinetic [KN75]. kinetics [BH77]. knowledge [Jør78b]. Kootenay [Clo78]. Kraus [Nih78].

L [Jør77e]. L. [RM79]. Lake [CKDF79, Clo78, HUS79, HM78, IA76, IA78, IG78, Jør76f, JMF78, Kal78, LN78, HL78, IA78, KN75, Mit76]. lakes [DLvS78, HL78, Kam78, NW79, Nyh78]. lands [Kri78]. Langmuir [Led79]. Large [PFF76, HL78]. layers [Nih78]. lead [WRR78]. lecturers [Jes76]. Lepomis [CBB77]. Lessons [CJH79]. limit [Goh77]. limited [Goh77]. limnology [Jør77e]. linear [Agn79, Cal79, Gop77, Hal76, Hal77, HR79]. litter [DW78, LBD78]. loading [BD78]. London [Jør79b, Jør79a]. long [BBMT75, Hil79, Jør76d]. long-term [BBMT75]. Lotka [GJ77]. Ltd [Jør77b].

M [Jac79]. macrochirus [CBB77]. making [Bis75a, Bis78]. Maloney [DM75]. mammal [Fra79]. man [Jør79c, Zuc75]. management [Bis75b, CJH79, Mit76, Pet77, VLG77, Jør77a, Rey77]. Manten [Jes76]. Marcel [Jør75b]. March [Ano76e]. Marine [Nih78, Jør75a]. Markovian [Bar78a, Bar78b, Bar79]. marsh [BO76]. Martin [Jac79]. mass [FKBB75]. material [Swa79]. Mathematical [Bis75a, Kal78, WVW78, CBB77, HB78, Jør75b, Mej79, Nih76, SLM^{+}79, Sei79, Zuc75, Jac79, Mej79]. mathematics [Cal79]. May [Ano75d, Ano77m, Ano78m, Ano79o]. Maynard [Jør75d]. McGraw [Rey77]. McGraw-Hill [Rey77]. measurable [SE76]. mechanisms [Str79]. meetings [Jes76]. mercury [CBB77]. Metabolism [McK77]. metal [GWJ79, Jør76e, WGWJ79]. metals [Jør79e, Sei79, Con75]. methodology [WM76]. Methods [Swa79]. methyl [CBB77]. MI [Gro79, Jac79]. Miami [DL77, Zuc75]. Mich [Con75, DM75]. Microclimate [Els76]. microcosm [BH77]. Middlebrooks [DM75]. migration [Kam78]. mixed [AA78]. Model [HUS79, AAV78, Bar78a, Bar78b, BD78, BV78b, Clo78, Cri78, DL77, DLB78a, DLB78b, DW78, Dub75, EWS78, Gil77, GWJ79, Hal77, HB78, IA78, IG78, JBB^{+}75, Jør75e, Jør76e, Jør76f, Jør76g, JMF78, Kal78, KM79, LN78, MP77, McK77, MKU79, Nih76, Nyh78, Par78b, Per78, RM79, SP76, SLM^{+}79, Sei79, Sjö77, WGWJ79, WRR78, YU79, DLB78a, DLB78b, WGWJ79, Sko79]. Modeling [Cal79, Jac79, Mit76, PLM75, Swa79, UFHH78]. Modelling [BH77, HM78, JM76a, JM76b, Jør79f, Jør79e, LBD78, BBMT75, BS78, Bis75a, Bis78, CHPS78, Get78, HUS79, Jør75c, Jør78a, Jør78b, Jør79d, Jør79b, JM79, Jør79g, Kin78, Mau79, Nih79, OFA75, PFF76, Ano77d, DM75, Gro79, Jør75a, Jør77b, Nih78, Jør76b]. Models [Abe75, Bar79, Ben79, CBB77, Day77, GJ77, Hal76, HR78, HS77, HR79, Hil79, Mag75, OR79, PP79, Par78a, Sha76, Str79, THR^{+}78, WVW78, Zuc75, Jør75d]. Monographs [Jør75b, Jac75]. Montreal [Jør77b]. morphology [NW79]. mortality [Fra79]. movement [WRR78]. mud [JKNJ75]. mud-water [JKNJ75]. Multiobjective [Jør76c]. multiple [PFF76].

N [Bis78, Els76, Jør75b, Nih79]. N.S.W [Nih78]. N.Y [Els76, Nih78, Rey77, Smi77]. NATO [Nih78]. Natural [Str79, Jør79c, SW75]. nature [Zuc75]. neo [Jør79c]. neo-technic [Jør79c]. Nihoul [Jør75a]. nitrogen [IA76, JJ75, JM76a]. non [Agn79, NW79]. non-linear [Agn79]. non-stratified [NW79]. normally [Fra79]. North [Nih76]. Norwegian [SLM^{+}79]. Note [Ano77i, Bar79]. November [Ano78l]. Numbers [Ren79, Fra79]. Nutrient [Par78a, DLB78a, DLB78b, Nyh78, Par78b]. nutrient-plankton [Par78b]. nutrification [YU79].

observed [Fra79]. ocean [DL77, Nih78]. Oceanography [Jør75a]. October [Ano77l, Ano79n]. oil [SW75]. Ont [Jør77b, Nih78]. Optimal [Agn79]. optimization [Jør76c]. Optimum [Goh77, HBJ76]. options [Pet77]. order [HUS79]. organic [MKU79]. Organism [Kam78]. organizations [Les76]. organizers [Jes76]. outfall [DL77]. ovata [Clo78]. Oxford [Jør79c, Mid79, Mor79, Nih78]. oxygen [WVW78].

P [Gro79, Jac79, BD78, HUS79]. Pa [Jør77d]. Pages [Ano75f, Ano75g, Ano75d, Ano75e, Ano76g, Ano76h, Ano76e, Ano76f, Ano77k, Ano77j, Ano77m, Ano77l, Ano78n, Ano78i, Ano78j, Ano78o, Ano78m, Ano78l, Ano78k, Ano79p, Ano79k, Ano79l, Ano79q, Ano79o, Ano79n, Ano79m]. Papers [Jør77d]. paradox [KM79]. Parameter [Ben79, LN78, SP76]. parameterization [Bar79]. Paravespula [RM79]. part [Gil77, Ren79]. parts [Ren79]. patchiness [Dub75, Kam78, Led79]. Patten [Jør79a]. patterns [Get78, Par78b]. peat [BO76]. pelagic [Sjö77]. Pergamon [Nih78, Nih78]. periodicity [Par75]. persistence [Goh75]. pest [Pet77]. phenol [BV78a]. phosphate [JKNJ75]. phosphorus [Arn78, BO76, JM76a, KN75, LN78, MKU79]. Physiological [Jør77e]. phytoplankton [Arn78, KM79, Nyh78]. piecewise [Cal79]. Pielou [Mej79, Smi77]. plankton [Dub75, KM79, Led79, Par78b]. planktonic [Kam78]. planning [Abe75, Jør79c]. Plant [Gil77, HJ76, McK77]. plume [DL77]. policy [CJH79, Jør76b, PLM75]. pollution [Jac79, Jør75b]. polychlorinated [KK79]. polyhedral [CKDF79]. population [Clo78, Fra79, Gop77, Jør75e, Kal78, RM79]. populations [Dub75]. power [McK77]. pp [Bis78, Con75, DM75, Els76, Gro79, Jac75, Jac79, Jes76, Jør75d, Jør75a, Jør75b, Jør76a, Jør76c, Jør77a, Jør77d, Jør77b, Jør77e, Jør79b, Jør79c, Jør79a, Mej79, Mid79, Mor79, Nih78, Nih79, Ren79, Rey77, Sko79, Smi77]. practice [Jør79b]. Praeger [Bis78]. Pratt [Jør75b]. predator [AAV78, Dub75, Hil79]. predicting [DLB78a, DLB78b]. prediction [IG78, Mag75, Nih78, PP79]. present [CHPS78]. Press [Jør75d, Jør79a, Nih78]. prey [AAV78, Dub75, Hil79, Jac75]. price [Els76]. primary [Cal79]. Problems [Bis78, BB75, Kin78, Jør75b, Nih79]. Procedures [GR77, LN78]. Proceedings [Nih78, Ren79]. process [JBB^{+}75, SE76, GWJ79, DM75]. processes [Gro79, HL78, Jac79]. production [DG79, Jør75e]. productivity [Cal79]. programming [HBJ76]. proposed [WM76]. publications [Jes76]. publisher [Ano77i]. Publishers [Bis78, Con75, DM75, Gro79, Jac79]. Publishing [Jes76, Jør77a, Rey77].

Pudoc [Jac75].

quality [Abe75, Bis75b, CHPS78, IA78, Mau79, OFA75, PFF76, WM76].

randomly [GJ77]. rate [Fra79]. Raymond [Gro79]. Recreation [Get78]. recycling [Par78a, SW75]. region [Zuc75]. Relating [PLM75]. Relative [Hal76]. release [JJ75]. representation [CKDF79]. reproduction [Fra79]. Research [Bis78, Foe76, Jør77a]. resource [Bar78b]. resources [Jør76a, Jør76c, Jør77b, SW75]. response [Jac75, Mag75]. resulting [Kam78]. Results [HB78, Jør78b, OFA75]. retrospective [Bar78b]. Review [KK79]. Revised [Mej79]. Richard [Jør77d]. river [BV78a]. role [Mau79]. root [VA76]. Rosenberg [Els76]. Ross [Jør77d]. Rubin [Con75].

S [Jør79b, Ren79, Sko79]. salt [HM78]. San [Jør79a]. satiation [AAV78]. scale [PFF76]. Science [Con75, DM75, Gro79, Jac79, Mid79, Jør77e, Jør76a, Jør76c]. Scientific [Bis78, Jes76]. Sea [HB78, Nih76]. season [Goh77]. sediment [KN75]. sediment-water [KN75]. sediments [JJ75]. seiche [Kam78]. selected [SP76]. selectivity [VS79]. semi [Kri78]. semi-arid [Kri78]. Seminar [Bis75b]. Sensitivity [Cri78, Hal77]. September [Ano75e, Ano76f, Ano78k, Ano79m, BB75]. Series [Jør75a, Mid79, Ren79, EWS78, Mor79, Nih78]. shallow [DLvS78, HM78, Nyh78]. Simulating [DG79]. Simulation [BO76, Clo78, Phi79, Ren79, Swa79, IA78, JBB^{+}75, Jør76d, Jør79a, Kal78, Mid79, MKU79, Nyh78, Sko79, VA76, WHP75, WRR78, Jac75, Nih79]. simulators [HJ76]. size [Boy79, Goh77]. small [Cri78, HUS79]. Smith [Jør75d]. smoothing [DG79]. soil [BV78b, DW78, MKU79]. solute [LBD78]. solution [Hal77]. Some [BBMT75, HJ76, Hil79, THR^{+}78, Kri78]. Sons [Smi77]. southern [Clo78, Nih76]. Spatial [Par78b, MP77]. special [Gil77]. species [BS78, Per78]. Spectral [EWS78]. Springer [Sko79]. Springer-Verlag [Sko79]. Stability [Goh75, GJ77, Hal76, IA76, Par75]. stand [DLB78a, DLB78b]. State [Ano76i, IG78]. State-of-the-art [Ano76i]. stationary [Cal79]. Statistical [Ano77h, BH77, Jør75b]. Statistics [Jør75b]. Stochastic [Day77, PP79, BD78, THR^{+}78]. strategies [MP77]. strategy [Per78]. Strathclyde [Mid79]. stratified [NW79]. streams [Cri78, PP79]. stressed [Get78]. Stroudsburg [Jør77d]. structure [HS77, Mag75]. studies [Sko79]. Study [BV78a, Nih78, CJH79, IA78, LN78, Sjö77, Zuc75]. submodel [Gil77, JJ75, JKNJ75]. subsidence [BV78b]. substances [Jør79f]. suggestions [Jør76d]. summer [NW79]. sunfish [CBB77]. Superior [HL78]. Sydney [Jør79b, Nih78]. Symposia [Jes76]. symposium [Jes76]. system [Arn78, Hal77, HUS79, Pet77, Jac75]. Systems [BB75, BV78b, Foe76, Les76, GR77, Jør76a, Jør76c, Jør77a, Jør77b, PFF76, Sha76, Sjö77, VLG77, Zuc75, Jør75a, Ren79, Jør79a, Rey77].

T [DM75, Jør76c, Jør77a]. tables [Jør76a, Jør76c, Jør77d, Jør77e, Mor79, Nih78]. targets [VLG77]. technic [Jør79c]. technique [Kin78]. techniques [Mid79]. temperatures [Cri78]. temporal [Get78]. term [BBMT75, Hil79, Jør76d]. terrestrial [GWJ79, WGWJ79]. Textbooks [Jør75b]. their [JM76a]. theoretical [Cri78]. theory [HUS79, Jør79b, WGWJ79, Zuc75]. Thermodynamic [BS78]. Thomas [Jac79]. time [Bar79, EWS78]. Toronto [Jør79b, Nih78]. toxic [Jør79f, KK79]. trace [DLB78a, DLB78b]. transformation [MKU79]. transformations [BV78b]. Transport [KK79, FKBB75, GWJ79, WGWJ79]. trophic [IG78]. trophogenic [HUS79]. tundra [DW78]. Turbulence [KM79].

U.S [Gro79, Jac79, Jør79c, Mid79, Sko79]. U.S.A [Gro79, Jac79]. University [Jør75d, Mid79]. unstable [Sjö77]. upper [Nih78]. uptake [CBB77, LBD78, Sei79]. urban [Zuc75]. Urbino [Nih78]. use [GR77, SE76, VS79]. using [Cal79, DG79].

V [Mor79]. value [OFA75]. values [SP76]. various [CBB77]. vegetation [LBD78, Phi79]. Verlag [Sko79]. vertical [Kam78]. via [DW78]. view [AA78]. viii [Gro79]. Vol [Jør75a, Jør75b, Jør79a]. Volterra [GJ77]. Volume [Jør76a, Jør76c, Jør77d, Jør77e, Mor79, Nih78, Ren79, Sko79]. vulgaris [RM79]. vulnerability [Goh75].

W [Jør75b, Jør76c, Jør77a, Jør79b, Nih79]. Wageningen [Jac75, Jør77a]. Wanielista [Jac79]. wasp [RM79]. Water [Jør77e, Abe75, BO76, Bis75b, BV78a, CHPS78, Cri78, DLB78a, DLB78b, HM78, IA78, Jac79, JKNJ75, Jør76a, Jør76e, Jør76c, Jør77b, KN75, Mau79, OFA75, PFF76, Rey77, Sha76, VA76, WVW78]. waterhyacinth [Mit76]. watershed [WRR78]. weevil [JBB^{+}75]. well [AA78]. well-mixed [AA78]. Wet [Mor79]. wetlands [Phi79]. Wiegert [Jør77d]. wildlife [WM76]. Wiley [DM75, Els76, Jør79b, Mej79, Nih79, Smi77]. Wiley-Interscience [Els76]. Wind [DLvS78]. Wit [Jør77a]. within [HUS79]. workshop [Ano76i]. World [Mor79, Jør75e].

X [Jør76a]. xv [Nih79].

York [Bis78, Els76, Jør75b, Jør79b, Jør79c, Jør79a, Mej79, Mid79, Mor79, Nih78, Rey77, Sko79, Smi77, OFA75, YU79]. Yugoslavia [BB75].

zone [HUS79, VA76]. Zooplankton [VS79, Arn78, Kal78, SE76].

References

Aydin:1978:NVD

[AA78] Fehmi N. Aydin and Robert C. Ahlert. A new view of dispersion in well-mixed estuaries. Ecological Modelling, 5(4):301–326, November 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900406.

Arditi:1978:PPM

[AAV78] R. Arditi, J. M. Abillon, and J. Vieira Da Silva. A predator– prey model with satiation and intraspecific competition. Ecological Modelling, 5(3):173–191, September 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900194.

Abendt:1975:MWQ

[Abe75] R. W. Abendt. Models in water quality planning. Ecological Modelling, 1(3):205–217, September 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380075900186.

Agnew:1979:OEF

[Agn79] T. T. Agnew. Optimal exploitation of a fishery employing a non-linear harvesting function. Ecological Modelling, 6(1):47–57, January 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900577.

Anonymous:1975:AI

[Ano75a] Anonymous. Author index. Ecological Modelling, 1(4):319–320, December 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900150.

Anonymous:1975:EB

[Ano75b] Anonymous. Editorial Board. Ecological Modelling, 1(1):iii, May 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380075900010.

Anonymous:1975:FE

[Ano75c] Anonymous. Forthcoming event. Ecological Modelling, 1(3):240, September 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900241.

Anonymous:1975:PM

[Ano75d] Anonymous. Pages 1–89 (May 1975). Ecological Modelling, 1(1):??, May 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1975:PS

[Ano75e] Anonymous. Pages 163–240 (September 1975). Ecological Modelling, 1(3):??, September 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1975:PD

[Ano75f] Anonymous. Pages 241–320 (December 1975). Ecological Modelling, 1(4):??, December 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1975:PJ

[Ano75g] Anonymous. Pages 91–162 (July 1975). Ecological Modelling, 1 (2):??, July 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1975:RB

[Ano75h] Anonymous. Recent books. Ecological Modelling, 1(1):83–84, May 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380075900095.

Anonymous:1976:A

[Ano76a] Anonymous. Announcement. Ecological Modelling, 2(3):250, September 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900260.

Anonymous:1976:AI

[Ano76b] Anonymous. Author index. Ecological Modelling, 2(4):327–328, December 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007690020X.

Anonymous:1976:EB

[Ano76c] Anonymous. Editorial Board. Ecological Modelling, 2(1):iii, March 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900016.

Anonymous:1976:E

[Ano76d] Anonymous. Erratum. Ecological Modelling, 2(4):325, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380076900193.

Anonymous:1976:PM

[Ano76e] Anonymous. Pages 1–89 (March 1976). Ecological Modelling, 2 (1):??, March 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1976:PS

[Ano76f] Anonymous. Pages 169–250 (September 1976). Ecological Modelling, 2(3):??, September 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1976:PD

[Ano76g] Anonymous. Pages 251–328 (December 1976). Ecological Modelling, 2(4):??, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1976:PJ

[Ano76h] Anonymous. Pages 91–167 (June 1976). Ecological Modelling, 2 (2):??, June 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1976:SAW

[Ano76i] Anonymous. State-of-the-art workshop. Ecological Modelling, 2 (4):325, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380076900181.

Anonymous:1977:Aa

[Ano77a] Anonymous. Announcement. Ecological Modelling, 3(1):77, February 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380077900266.

Anonymous:1977:Ab

[Ano77b] Anonymous. Announcement. Ecological Modelling, 3(2):165, May 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380077900059.

Anonymous:1977:AI

[Ano77c] Anonymous. Author index. Ecological Modelling, 3(4):321–322, October 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900187.

Anonymous:1977:CEM

[Ano77d] Anonymous. Conference on ecological modelling. Ecological Modelling, 3(4):319, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380077900151.

Anonymous:1977:EB

[Ano77e] Anonymous. Editorial Board. Ecological Modelling, 3(1):iii, February 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900199.

Anonymous:1977:Ea

[Ano77f] Anonymous. Erratum. Ecological Modelling, 3(2):166, May 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380077900060.

Anonymous:1977:Ec

[Ano77g] Anonymous. Erratum. Ecological Modelling, 3(3):248, August 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380077900333.

Anonymous:1977:FIA

[Ano77h] Anonymous. Formation of the International Association for Statistical Computing. Ecological Modelling, 3(4):320, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380077900163.

Anonymous:1977:NP

[Ano77i] Anonymous. Note from the publisher. Ecological Modelling, 3(4): 320, October 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900175.

Anonymous:1977:PF

[Ano77j] Anonymous. Pages 1–84 (February 1977). Ecological Modelling, 3 (1):??, February 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1977:PA

[Ano77k] Anonymous. Pages 167–248 (August 1977). Ecological Modelling, 3 (3):??, August 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1977:PO

[Ano77l] Anonymous. Pages 249–322 (October 1977). Ecological Modelling, 3 (4):??, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1977:PM

[Ano77m] Anonymous. Pages 85–166 (May 1977). Ecological Modelling, 3 (2):??, May 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1978:Aa

[Ano78a] Anonymous. Announcement. Ecological Modelling, 4(2–3):311, February 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900121.

Anonymous:1978:Ac

[Ano78b] Anonymous. Announcement. Ecological Modelling, 5(3):258, September 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900248.

Anonymous:1978:Ab

[Ano78c] Anonymous. Announcements. Ecological Modelling, 4(4):371– 372, May 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900297.

Anonymous:1978:AIa

[Ano78d] Anonymous. Author index. Ecological Modelling, 4(4):373– 374, May 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900315.

Anonymous:1978:AIb

[Ano78e] Anonymous. Author index. Ecological Modelling, 5(4):327–328, November 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900418.

Anonymous:1978:EBa

[Ano78f] Anonymous. Editorial Board. Ecological Modelling, 4(1):iii, January 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007890042X.

Anonymous:1978:EBb

[Ano78g] Anonymous. Editorial Board. Ecological Modelling, 5(1):iii, June 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900133.

Anonymous:1978:E

[Ano78h] Anonymous. Erratum. Ecological Modelling, 4(4):372, May 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900303.

Anonymous:1978:PJa

[Ano78i] Anonymous. Pages 1–73 (January 1978). Ecological Modelling, 4 (1):??, January 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1978:PJb

[Ano78j] Anonymous. Pages 1–90 (June 1978). Ecological Modelling, 5(1):??, June 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1978:PS

[Ano78k] Anonymous. Pages 173–258 (September 1978). Ecological Modelling, 5(3):??, September 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1978:PN

[Ano78l] Anonymous. Pages 259–328 (November 1978). Ecological Modelling, 5(4):??, November 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1978:PM

[Ano78m] Anonymous. Pages 313–374 (May 1978). Ecological Modelling, 4 (4):??, May 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1978:PF

[Ano78n] Anonymous. Pages 77–311 (February 1978). Ecological Modelling, 4(2–3):??, February 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1978:PJc

[Ano78o] Anonymous. Pages 91–171 (July 1978). Ecological Modelling, 5 (2):??, July 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1979:Aa

[Ano79a] Anonymous. Announcement. Ecological Modelling, 6(1):88, January 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900607.

Anonymous:1979:Ab

[Ano79b] Anonymous. Announcement. Ecological Modelling, 6(2):175–177, February 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007990036X.

Anonymous:1979:Ad

[Ano79c] Anonymous. Announcement. Ecological Modelling, 7(1):83, June 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900103.

Anonymous:1979:Ae

[Ano79d] Anonymous. Announcement. Ecological Modelling, 7(3):238, September 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900723.

Anonymous:1979:Ac

[Ano79e] Anonymous. Announcements. Ecological Modelling, 6(4):337– 338, May 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900504.

Anonymous:1979:AIa

[Ano79f] Anonymous. Author index. Ecological Modelling, 6(4):341– 342, May 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900528.

Anonymous:1979:AIb

[Ano79g] Anonymous. Author index. Ecological Modelling, 7(4):313, October 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900413.

Anonymous:1979:EBa

[Ano79h] Anonymous. Editorial Board. Ecological Modelling, 6(1):iii, January 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007990053X.

Anonymous:1979:EBb

[Ano79i] Anonymous. Editorial Board. Ecological Modelling, 7(1):iii, June 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900073.

Anonymous:1979:E

[Ano79j] Anonymous. Errata. Ecological Modelling, 6(4):339, May 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900516.

Anonymous:1979:PJa

[Ano79k] Anonymous. Pages 1–88 (January 1979). Ecological Modelling, 6 (1):??, January 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1979:PJb

[Ano79l] Anonymous. Pages 1–90 (June 1979). Ecological Modelling, 7(1):??, June 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1979:PS

[Ano79m] Anonymous. Pages 165–238 (September 1979). Ecological Modelling, 7(3):??, September 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1979:PO

[Ano79n] Anonymous. Pages 239–313 (October 1979). Ecological Modelling, 7 (4):??, October 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1979:PM

[Ano79o] Anonymous. Pages 289–342 (May 1979). Ecological Modelling, 6 (4):??, May 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1979:PF

[Ano79p] Anonymous. Pages 89–177 (February 1979). Ecological Modelling, 6 (2):??, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1979:PJc

[Ano79q] Anonymous. Pages 91–163 (July 1979). Ecological Modelling, 7 (2):??, July 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Arnold:1978:AZP

[Arn78] E. M. Arnold. Aspects of a zooplankton, phytoplankton, phosphorus system. Ecological Modelling, 5(4):293–300, November 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030438007890039X.

Barber:1978:MME

[Bar78a] Mahlon Craig Barber. A Markovian model for ecosystem flow analysis. Ecological Modelling, 5(3):193–206, September 1978. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380078900200.

Barber:1978:RMM

[Bar78b] Mahlon Craig Barber. A retrospective Markovian model for ecosystem resource flow. Ecological Modelling, 5(2):125–135, July 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900352.

Barber:1979:NCT

[Bar79] Mahlon Craig Barber. A note concerning time parameterization of Markovian models of ecosystem flow analysis. Ecological Modelling, 6(4):323–328, May 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380079900449.

Biswas:1975:SAE

[BB75] Margaret R. Biswas and Asit K. Biswas. Systems approach to environmental problems, September 8th–17th, 1975, Dubrovnik, Yugoslavia. Ecological Modelling, 1(3):239–240, September 1975. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030438007590023X.

Behrens:1975:SAM

[BBMT75] Jens C. Behrens, Jan E. Beyer, Oli B. G. Madsen, and Per G. Thomsen. Some aspects of modelling the long-term behaviour of aquatic ecosystems. Ecological Modelling, 1(3):163–198, September 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380075900162.

Bogardi:1978:ISC

[BD78] István Bogárdi and Lucien Duckstein. Input for a stochastic control model of P loading. Ecological Modelling, 4(2–3):173–195, February 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900066.

Benson:1979:PFD

[Ben79] M. Benson. Parameter fitting in dynamic models. Ecological Modelling, 6(2):97–115, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900292.

Bargmann:1977:EAS

[BH77] Rolf E. Bargmann and Efraim Halfon. Efficient algorithms for statistical estimation in compartmental analysis: Modelling ^{60}Co kinetics in an aquatic microcosm. Ecological Modelling, 3(3):211– 226, August 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900291.

Biswas:1975:MME

[Bis75a] Asit K. Biswas. Mathematical modelling and environmental decision-making. Ecological Modelling, 1(1):31–48, May 1975. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380075900046.

Biswas:1975:SWQ

[Bis75b] Asit K. Biswas. Seminar on water quality management. Ecological Modelling, 1(2):161–162, July 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380075900319.

Biswas:1978:EMD

[Bis78] Asit K. Biswas. Environmental modelling and decision making: the Holcomb Research Institute for the Scientific Committee on Problems of the Environment. Praeger Publishers, Inc., New York, N. Y., 1976, 152 pp., US $16.50, ISBN 0-275-24190-4. Ecological Modelling, 4(1):71–72, January 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900509.

Bayley:1976:SIE

[BO76] Suzanne Bayley and Howard T. Odum. Simulation of interrelations of the Everglades’ marsh, peat, fire, water and phosphorus. Ecological Modelling, 2(3):169–188, September 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380076900211.

Boyd:1979:SAH

[Boy79] A. V. Boyd. The size of animal herds. Ecological Modelling, 6 (2):91–96, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380079900280.

Bhagat:1978:TMB

[BS78] Phiroz M. Bhagat and Richard E. Sonntag. Thermodynamic modelling of bio-species accommodation. Ecological Modelling, 4(1):61– 66, January 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900480.

Borighem:1978:SBP

[BV78a] G. Borighem and J. Vereecken. Study of the biodegradation of phenol in river water. Ecological Modelling, 4(1):51–59, January 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900479.

Browder:1978:SMC

[BV78b] Joan A. Browder and Bob G. Volk. Systems model of carbon transformations in soil subsidence. Ecological Modelling, 5(4):269–292, November 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900388.

Cale:1979:MGP

[Cal79] William G. Cale. Modeling grassland primary productivity using piecewise stationary, piecewise linear mathematics. Ecological Modelling, 7(2):107–123, July 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900024.

Curtis:1977:AVM

[CBB77] E. H. Curtis, J. J. Beauchamp, and B. G. Blaylock. Application of various mathematical models to data from the uptake of methyl mercury in bluegill sunfish (Lepomis macrochirus). Ecological Modelling, 3(4):273–284, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380077900084.

Cembrowicz:1978:APH

[CHPS78] R. G. Cembrowicz, H. H. Hahn, E. J. Plate, and G. A. Schultz. Aspects of present hydrological and water quality modelling. Ecological Modelling, 5(1):39–66, June 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900169.

Clark:1979:LEP

[CJH79] William C. Clark, Dixon D. Jones, and C. S. Holling. Lessons for ecological policy design: a case study of ecosystem management. Ecological Modelling, 7(1):1–53, June 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900085.

Casti:1979:LEP

[CKDF79] J. Casti, J. Kempf, L. Duckstein, and M. Fogel. Lake ecosystems: a polyhedral dynamics representation. Ecological Modelling, 7(3): 223–237, September 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380079900711.

Cloern:1978:SMC

[Clo78] James E. Cloern. Simulation model of Cryptomonas ovata population dynamics in southern Kootenay Lake, British Columbia. Ecological Modelling, 4(2–3):133–149, February 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380078900042.

Connell:1975:AEC

[Con75] D. W. Connell. Aqueous-environmental chemistry of metals: Alan J. Rubin (Editor). Ann Arbor Science Publishers Inc., Ann Arbor, Mich., 1974, 390 pp., £9.80. Ecological Modelling, 1(3):235–236, September 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900204.

Crittenden:1978:SAT

[Cri78] Robert N. Crittenden. Sensitivity analysis of a theoretical energy balance model for water temperatures is small streams. Ecological Modelling, 5(3):207–224, September 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900212.

Dayananda:1977:SMF

[Day77] P. W. A. Dayananda. Stochastic models for forest fires. Ecological Modelling, 3(4):309–313, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380077900114.

Dyer:1979:SFP

[DG79] T. G. J. Dyer and Jane F. Gilloly. Simulating fish production using exponential smoothing. Ecological Modelling, 6(1):77– 87, January 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900590.

DAmato:1977:KMC

[DL77] R. D’Amato and T. N. Lee. A kinematic model of the city of Miami ocean outfall plume behavior. Ecological Modelling, 3(3):227– 243, August 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900308.

Dixon:1978:PCMa

[DLB78a] K. R. Dixon, R. J. Luxmoore, and C. L. Begovich. CERES — a model of forest stand biomass dynamics for predicting trace contaminant, nutrient, and water effects. I. Model description. Ecological Modelling, 5(1):17–38, June 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380078900157.

Dixon:1978:PCMb

[DLB78b] K. R. Dixon, R. J. Luxmoore, and C. L. Begovich. CERES — a model of forest stand biomass dynamics for predicting trace contaminant, nutrient, and water effects. II. Model application. Ecological Modelling, 5(2):93–114, July 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380078900339.

DiGiano:1978:WID

[DLvS78] Francis A. DiGiano, Lambertus Lijklema, and Gerrit van Straten. Wind induced dispersion and algal growth in shallow lakes. Ecological Modelling, 4(2–3):237–252, February 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900091.

Dahl-Madsen:1975:MEP

[DM75] K. I. Dahl-Madsen. Modelling the eutrophication process: E. J. Middlebrooks, D. H. Falkenborg and T. E. Maloney (Editors). Ann Arbor Science Publishers, Ann Arbor, Mich. (distributed in Great Britain by Wiley, Chichester), 1974, 228 pp., £7.90 ($14.50). Ecological Modelling, 1(1):81–82, May 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380075900071.

Dubois:1975:MPP

[Dub75] Daniel M. Dubois. A model of patchiness for prey — predator plankton populations. Ecological Modelling, 1(1):67–80, May 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030438007590006X.

Douce:1978:IES

[DW78] G. Keith Douce and David P. Webb. Indirect effects of soil invertebrates on litter decomposition: Elaboration via analysis of a tundra model. Ecological Modelling, 4(4):339–359, May 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900273.

Elston:1976:MBE

[Els76] J. Elston. Microclimate: the biological environment: N. J. Rosenberg. Wiley-Interscience, New York, N.Y., 1974, 315 pp., price £6.40. Ecological Modelling, 2(3):249, September 1976. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380076900259.

Emanuel:1978:SAF

[EWS78] W. R. Emanuel, D. C. West, and H. H. Shugart. Spectral analysis of forest model time series. Ecological Modelling, 4(4):313– 326, May 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007890025X.

Falco:1975:EMT

[FKBB75] James W. Falco, Pat C. Kerr, Michael B. Barron, and Donald L. Brockway. The effect of mass transport on biostimulation of algal growth. Ecological Modelling, 1(2):117–131, July 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380075900277.

Foell:1976:EER

[Foe76] W. K. Foell. Ecological and environmental research at the International Institute for Applied Systems Analysis. Ecological Modelling, 2(4):287–294, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380076900119.

Fransz:1979:EBR

[Fra79] H. G. Fransz. Estimation of birth rate and juvenile mortality from observed numbers of juveniles in a mammal population with normally dispersed reproduction. Ecological Modelling, 7(2):125– 133, July 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900036.

Getz:1978:MTP

[Get78] Wayne M. Getz. On modelling temporal patterns in stressed ecosystems — recreation in a coniferous forest. Ecological Modelling, 5 (3):237–257, September 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380078900236.

Gilmanov:1977:PSH

[Gil77] T. G. Gilmanov. Plant submodel in the holistic model of a grassland ecosystem (with special attention to the belowground part). Ecological Modelling, 3(2):149–163, May 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380077900047.

Goh:1977:FSR

[GJ77] B. S. Goh and L. S. Jennings. Feasibility and stability in randomly assembled Lotka–Volterra models. Ecological Modelling, 3 (1):63–71, February 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380077900242.

Goh:1975:SVP

[Goh75] B. S. Goh. Stability, vulnerability and persistence of complex ecosystems. Ecological Modelling, 1(2):105–116, July 1975. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380075900265.

Goh:1977:OSL

[Goh77] B. S. Goh. Optimum size limit for a fishery with a limited fishing season. Ecological Modelling, 3(1):3–15, February 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380077900217.

Gopalsamy:1977:ADP

[Gop77] K. Gopalsamy. Age dependent population dispersion in linear habitats. Ecological Modelling, 3(2):119–132, May 1977. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380077900023.

Gilliland:1977:USD

[GR77] Martha W. Gilliland and Paul G. Risser. The use of systems diagrams for environmental impact assessment: Procedures and an application. Ecological Modelling, 3(3):183–209, August 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030438007790028X.

Gromiec:1979:MBP

[Gro79] M. J. Gromiec. Modelling biochemical processes in aquatic ecosystems: Raymond P. Canale (Editor). Ann Arbor Science Publishers, Ann Arbor, MI 48106, U.S.A., 1976, viii + 389 pp., £15.55, U.S. $24.75, ISBN: 250-401045. Ecological Modelling, 6(4):329– 330, May 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900450.

Grenney:1979:ATH

[GWJ79] W. J. Grenney, R. J. Wagenet, and J. J. Jurinak. An atmospheric — terrestrial heavy metal transport model. II. Process equations. Ecological Modelling, 6(3):273–288, March 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900164.

Halfon:1976:RSE

[Hal76] Efraim Halfon. Relative stability of ecosystem linear models. Ecological Modelling, 2(4):279–286, December 1976. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380076900107.

Halfon:1977:ASS

[Hal77] Efraim Halfon. Analytical solution of the system sensitivity equation associated with a linear model. Ecological Modelling, 3(4): 301–307, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380077900102.

Horwood:1978:RHM

[HB78] J. W. Horwood and J. A. Bedwell. Results from a hydrodynamical mathematical model of the Irish Sea. Ecological Modelling, 4(4):327–337, May 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380078900261.

Hunter:1976:OFA

[HBJ76] Dennis H. Hunter, E. T. Bartlett, and Donald A. Jameson. Optimum forage allocation through chance-constrained programming. Ecological Modelling, 2(2):91–99, June 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380076900272.

Hilborn:1979:SLT

[Hil79] Ray Hilborn. Some long term dynamics of predator–prey models with diffusion. Ecological Modelling, 6(1):23–30, January 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900553.

Hesketh:1976:SCC

[HJ76] J. D. Hesketh and J. W. Jones. Some comments on computer simulators for plant growth — 1975. Ecological Modelling, 2(3): 235–247, September 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380076900247.

Halfon:1978:EAD

[HL78] E. Halfon and D. C. L. Lam. The effects of advection — diffusion processes on the eutrophication of large lakes, a hypothetical example: Lake Superior. Ecological Modelling, 4(2–3):119– 131, February 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900030.

Hutchison:1978:MWS

[HM78] I. P. G. Hutchison and D. C. Midgley. Modelling the water and salt balance in a shallow lake. Ecological Modelling, 4(2–3):211– 235, February 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007890008X.

Halfon:1978:AEM

[HR78] Efraim Halfon and Marcello G. Reggiani. Adequacy of ecosystem models. Ecological Modelling, 4(1):41–50, January 1978. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380078900467.

Harwell:1979:EAL

[HR79] Mark A. Harwell and Harvey L. Ragsdale. Eigengroup analyses of linear ecosystem models. Ecological Modelling, 7(4):239– 255, October 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900371.

Hallam:1977:ISD

[HS77] Thomas G. Hallam and Daniel S. Simberloff. On the intrinsic structure of differential equation models of ecosystems. Ecological Modelling, 3(3):167–182, August 1977. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380077900278.

Halfon:1979:MOE

[HUS79] E. Halfon, H. Unbehauen, and C. Schmid. Model order estimation and system identification theory and application to the modelling of ^{32}P kenetics within the trophogenic zone of a small lake. Ecological Modelling, 6(1):1–22, January 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900541.

Ikeda:1976:DNC

[IA76] Saburo Ikeda and Norihiko Adachi. Dynamics of the nitrogen cycle in a lake and its stability. Ecological Modelling, 2(3):213–234, September 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900235.

Ikeda:1978:DWQ

[IA78] Saburo Ikeda and Norihiko Adachi. A dynamic water quality model of Lake Biwa — a simulation study of the lake eutrophication. Ecological Modelling, 4(2–3):151–172, February 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900054.

Imboden:1978:DLM

[IG78] D. M. Imboden and R. Gächter. A dynamic lake model for trophic state prediction. Ecological Modelling, 4(2–3):77–98, February 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900017.

Jacobsen:1975:FRP

[Jac75] O. S. Jacobsen. The functional response to prey density in an acarine system: H. G. Franz. In: Simulation Monographs. Pudoc, Wageningen, 1974, 143 pp., Dfl. 20.00. Ecological Modelling, 1 (3):237, September 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380075900228.

Jacobsen:1979:MMW

[Jac79] O. S. Jacobsen. Mathematical modeling for water pollution control processes: Thomas M. Keinath and Martin P. Wanielista (Editors). Ann Arbor Science Publishers, Ann Arbor, MI 48106, U.S.A., 1975, 450 pp., 139 figs., £16.15, U.S. $25.85, ISBN 0-250-40077-4. Ecological Modelling, 6(2):173–174, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900358.

Jones:1975:BWC

[JBB^{+}75] J. W. Jones, H. D. Bowen, J. R. Bradley, R. E. Stinner, and R. S. Sowell. The boll weevil (Coleoptera: Curculionidae) feeding process: a simulation model. Ecological Modelling, 1(4):289– 302, December 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900125.

Jesse:1976:SSP

[Jes76] Andreas Jesse. Symposia and symposium publications — a guide for organizers, lecturers and editors of scientific meetings: Arie A. Manten. Elsevier Scientific Publishing company, Amsterdam, 1975, 164 pp., Dfl. 37.50, ISBN 0-444-41419-3. Ecological Modelling, 2 (4):323–324, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/030438007690017X.

Jacobsen:1975:SNR

[JJ75] Ole Stig Jacobsen and Sven Erik Jørgensen. A submodel for nitrogen release from sediments. Ecological Modelling, 1(2):147– 151, July 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900290.

Jorgensen:1975:SAM

[JKNJ75] Sven Erik Jørgensen, Lars Kamp-Nielsen, and Ole Stig Jacobsen. A submodel for anaerobic mud-water exchange of phosphate. Ecological Modelling, 1(2):133–146, July 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380075900289.

Jorgensen:1976:MGC

[JM76a] Sven Erik Jørgensen and Henning Mejer. Modelling the global cycle of carbon, nitrogen and phosphorus and their influence on the global heat balance. Ecological Modelling, 2(1):19–31, March 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900041.

Jorgensen:1976:MGH

[JM76b] Sven Erik Jørgensen and Henning Mejer. Modelling the global heat balance. Ecological Modelling, 2(4):273–277, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380076900090.

Jorgensen:1977:EBC

[JM77] Sven Erik Jørgensen and Henning Mejer. Ecological buffer capacity. Ecological Modelling, 3(1):39–45, 47, 49, 51, 53–61, February 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900230.

Jorgensen:1979:HAE

[JM79] S. E. Jørgensen and H. Mejer. A holistic approach to ecological modelling. Ecological Modelling, 7(3):169–189, September 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900681.

Jorgensen:1978:ELM

[JMF78] Sven Erik Jørgensen, Henning Mejer, and Mogens Friis. Examination of a lake model. Ecological Modelling, 4(2–3):253–278, February 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900108.

Jorgensen:1975:MMS

[Jør75a] S. E. Jørgensen. Modelling of marine systems: J. C. J. Nihoul (Editor). Elsevier Oceanography Series, Vol. 10. Elsevier, Amsterdam, 1974, 272 pp., Dfl. 70.00. Ecological Modelling, 1(3):236–237, September 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900216.

Jorgensen:1975:SMA

[Jør75b] S. E. Jørgensen. Statistical and mathematical aspects of pollution problems: John W. Pratt (Editor). Statistics: Textbooks and Monographs, Vol. 6. Marcel Dekker Inc., New York, N. Y., 1974, 392 pp., $19.50. Ecological Modelling, 1(4):317–318, December 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900149.

Jorgensen:1975:AEM

[Jør75c] Sven Erik Jørgensen. About “ecological modelling”. Ecological Modelling, 1(1):1–2, May 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380075900022.

Jorgensen:1975:MEJ

[Jør75d] Sven Erik Jørgensen. Models in ecology: J. Maynard Smith. Cambridge University Press, Cambridge, 1974, 146 pp., 48 figures, £3.70. Ecological Modelling, 1(1):82–83, May 1975. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380075900083.

Jorgensen:1975:WMG

[Jør75e] Sven Erik Jørgensen. A world model of the growth in production and population. Ecological Modelling, 1(3):199–203, September 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380075900174.

Jorgensen:1976:CSW

[Jør76a] Sven Erik Jørgensen. Computer systems and water resources. Developments in Water Science: Volume 1 G. Bugliarello and F. Gunther. Elsevier, Amsterdam, 1974, 202 pp., 16 tables, 38 illustrations, US $18.75, Dfl. 45.00, ISBN 0-444-41259-X. Ecological Modelling, 2 (4):321–322, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380076900156.

Jorgensen:1976:EME

[Jør76b] Sven Erik Jørgensen. “Ecological modelling” — editorial policy of the journal. Ecological Modelling, 2(1):1–2, March 1976. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380076900028.

Jorgensen:1976:MOW

[Jør76c] Sven Erik Jørgensen. Multiobjective optimization in water resources systems. Developments in water science: Volume 3 Y. Y. Haimes, W. A. Hall and H. T. Freedman. Elsevier, Amsterdam, 1975, 200 pp., 16 tables, 28 illustrations US $17.50, Dfl. 42.00, ISBN 0-444-41313-8. Ecological Modelling, 2(4):322–323, December 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900168.

Jorgensen:1976:CCL

[Jør76d] Sven Erik Jørgensen. COLTS-continuous long term simulation a call for suggestions. Ecological Modelling, 2(2):167, June 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380076900314.

Jorgensen:1976:EMH

[Jør76e] Sven Erik Jørgensen. An ecological model for heavy metal contamination of crops and ground water. Ecological Modelling, 2(1): 59–67, March 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900065.

Jorgensen:1976:EML

[Jør76f] Sven Erik Jørgensen. A eutrophication model for a lake. Ecological Modelling, 2(2):147–165, June 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380076900302.

Jorgensen:1976:MFG

[Jør76g] Sven Erik Jørgensen. A model of fish growth. Ecological Modelling, 2(4):303–313, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380076900132.

Jorgensen:1977:CES

[Jør77a] S. E. Jørgensen. Critical evaluation of systems analysis in ecosystems research and management: G. W. Arnold and C. T. de Wit (Editors). Centre for Agricultural Publishing and Documentation, Wageningen, 1976, 108 pp., 20 figures, Dfl. 20.00, ISBN 90-220-0593-3. Ecological Modelling, 3(3):247–248, August 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380077900321.

Jorgensen:1977:MWR

[Jør77b] S. E. Jørgensen. Modelling of water resources systems, I and II: A. K. Biswas (Editor). Harvest House Ltd., Montreal, Ont., 1974, 636 pp., 186 illustrations, US $20.00. Ecological Modelling, 3(3):245– 247, August 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007790031X.

Jorgensen:1977:AYJ

[Jør77c] Sven Erik Jørgensen. About your journal. Ecological Modelling, 3 (1):1, February 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900205.

Jorgensen:1977:EEV

[Jør77d] Sven Erik Jørgensen. Ecological energetics: Volume 4 of Benchmark Papers in Ecology. Richard G. Wiegert (Editor). Dowden, Hutchinson and Ross, Stroudsburg, Pa., 1976, 458 pp., 88 tables, 108 illustrations, £15.00, US $27.75, ISBN 0-470-94361-0. Ecological Modelling, 3(4):315–316, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380077900126.

Jorgensen:1977:PLV

[Jør77e] Sven Erik Jørgensen. Physiological limnology: Volume 2 of Developments in Water Science. H. L. Golterman. Elsevier, Amsterdam, 1975, 490 pp., 62 tables, 131 illustrations, US $47.95, Dfl. 125.00, ISBN 0-444-41270-0. Ecological Modelling, 3(4):317– 318, October 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007790014X.

Jorgensen:1978:FEM

[Jør78a] Sven Erik Jørgensen. Focus on ecological modelling. Ecological Modelling, 4(1):1–2, January 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900431.

Jorgensen:1978:IEK

[Jør78b] Sven Erik Jørgensen. Improvement of exchange of knowledge and results in the field of ecological modelling. Ecological Modelling, 5 (2):91–92, July 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900327.

Jorgensen:1979:SAS

[Jør79a] S. E. Jørgensen. Systems analysis and simulation in ecology, vol. IV: Bernard C. Patten (Editor). Academic Press, New York, San Francisco, London, 1976, 593 pp., £25.75, ISBN 0-12-547204-8. Ecological Modelling, 6(2):170–173, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900346.

Jorgensen:1979:EMT

[Jør79b] Sven Erik Jørgensen. Ecosystem modelling in theory and practice. An introduction with case histories: Charles A. S. Hall and John W. Day, Jr. (Editors). John Wiley, New York, London, Sydney, Toronto, 1977, 684 pp., $38.00. Ecological Modelling, 6(4): 332–334, May 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900474.

Jorgensen:1979:PNM

[Jør79c] Sven Erik Jørgensen. Planning for natural man in the neo-technic environment: an ethological approach: Barrie B. Greenbie. Elsevier, Amsterdam, Oxford, New York, 1976, 209 pp., 44 figures, Dfl. 75.00, U.S. $30.75, ISBN 0-444-41329-4. Ecological Modelling, 6(4):334–335, May 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380079900486.

Jorgensen:1979:CFE

[Jør79d] Sven Erik Jørgensen. Cooperation in the field of ecological modelling. Ecological Modelling, 6(2):89–90, February 1979. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380079900279.

Jorgensen:1979:MDEb

[Jør79e] Sven Erik Jørgensen. Modelling the distribution and effect of heavy metals in an aquatic ecosystem. Ecological Modelling, 6(3):199– 222, March 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900139.

Jorgensen:1979:MDEa

[Jør79f] Sven Erik Jørgensen. Modelling the distribution and effect of toxic substances. Ecological Modelling, 6(3):181, March 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900115.

Jorgensen:1979:NAE

[Jør79g] Sven Erik Jørgensen. New approaches in ecological modelling. Ecological Modelling, 7(3):167–168, September 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030438007990067X.

Kalmaz:1978:MMC

[Kal78] Ekrem V. Kalmaz. Mathematical model and computer simulation of the population dynamics of zooplankton in lake and estuary ecosystems. Ecological Modelling, 5(3):225–235, September 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900224.

Kamykowski:1978:OPL

[Kam78] Daniel Kamykowski. Organism patchiness in lakes resulting from the interaction between the internal seiche and planktonic diurnal vertical migration. Ecological Modelling, 4(2–3):197–210, February 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900078.

King:1978:ACM

[Kin78] G. A. M. King. Applying a chemical modelling technique to ecological problems. Ecological Modelling, 5(4):259–268, November 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900376.

Kalmaz:1979:TDT

[KK79] Ekrem V. Kalmaz and Gulgun D. Kalmaz. Transport, distribution and toxic effects of polychlorinated biphenyls in ecosystems: Review. Ecological Modelling, 6(3):223–251, March 1979. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380079900140.

Kemp:1979:TPD

[KM79] W. M. Kemp and W. J. Mitsch. Turbulence and phytoplankton diversity: a general model of the “paradox of plankton”. Ecological Modelling, 7(3):201–222, September 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438007990070X.

Kamp-Nielsen:1975:KAA

[KN75] Lars Kamp-Nielsen. A kinetic approach to the aerobic sedimentwater exchange of phosphorus in Lake Esrom. Ecological Modelling, 1(2):153–160, July 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380075900307.

Krishnamurthy:1978:HDS

[Kri78] L. Krishnamurthy. Herbage dynamics of some semi-arid Indian grazing lands. Ecological Modelling, 5(1):77–84, June 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900182.

Luxmoore:1978:MSU

[LBD78] R. J. Luxmoore, C. L. Begovich, and K. R. Dixon. Modelling solute uptake and incorporation into vegetation and litter. Ecological Modelling, 5(2):137–171, July 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900364.

Ledbetter:1979:LCP

[Led79] Max Ledbetter. Langmuir circulations and plankton patchiness. Ecological Modelling, 7(4):289–310, October 1979. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380079900395.

Lesourne:1976:SAI

[Les76] J. Lesourne. Systems analysis and international organizations. Ecological Modelling, 2(4):295–302, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380076900120.

Lewis:1978:SPE

[LN78] S. Lewis and A. Nir. A study of parameter estimation procedures of a model for lake phosphorus dynamics. Ecological Modelling, 4 (2–3):99–117, February 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380078900029.

Maguire:1975:APE

[Mag75] Bassett Maguire. Analysis and prediction of ecological and ecosystem dynamics with response structure models. Ecological Modelling, 1(4):269–287, December 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380075900113.

Mauersberger:1979:REW

[Mau79] Peter Mauersberger. On the role of entropy in water quality modelling. Ecological Modelling, 7(3):191–199, September 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900693.

McKellar:1977:MME

[McK77] Henry N. McKellar. Metabolism and model of an estuarine bay ecosystem affected by a coastal power plant. Ecological Modelling, 3(2):85–118, May 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380077900011.

Mejer:1979:MEC

[Mej79] Henning Mejer. Mathematical ecology: E. C. Pielou. (Revised edition of An introduction to mathematical ecology, 1966). Wiley, New York, 1977, x + 385 pp., US $22.45, £13.00, ISBN 0471-01993 3. Ecological Modelling, 6(2):169–170, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900334.

Midgley:1979:CST

[Mid79] D. C. Midgley. Computer simulation techniques in hydrology: George Fleming, University of Strathclyde. Environmental Science Series, Elsevier, New York, Oxford, Amsterdam, 1974, 333 pp., U.S. $33.95, Dfl. 85.00, ISBN 0-444-00157-3. Ecological Modelling, 6(4):331–332, May 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380079900462.

Mitsch:1976:EMW

[Mit76] William J. Mitsch. Ecosystem modeling of waterhyacinth management in Lake Alice, Florida. Ecological Modelling, 2(1):69– 89, March 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900077.

Mishra:1979:SMO

[MKU79] B. Mishra, P. K. Khanna, and B. Ulrich. A simulation model for organic phosphorus transformation in a forest soil ecosystem. Ecological Modelling, 6(1):31–46, January 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900565.

Morgan:1979:WCE

[Mor79] Neville C. Morgan. Wet coastal ecosystems: Volume 1 of the series Ecosystems of the World. V. J. Chapman (Editor). Elsevier, Amsterdam-Oxford-New York, 1977, 428 pp., 77 tables, 135 figures, US $49.95, Dfl. 122.00, ISBN 0-444-41560-2. Ecological Modelling, 6(2):167–168, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380079900322.

Maguire:1977:SMG

[MP77] Lynn A. Maguire and James W. Porter. A spatial model of growth and competition strategies in coral communities. Ecological Modelling, 3(4):249–271, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380077900072.

Nihoul:1976:IMM

[Nih76] Jacques C. J. Nihoul. An interdisciplinary mathematical model applied to the Southern Bight of the North Sea. Ecological Modelling, 2(1):3–17, March 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/030438007690003X.

Nihoul:1978:MPU

[Nih78] Jacques C. J. Nihoul. Modelling and prediction of the upper layers of the ocean: Proceedings of the NATO Advanced Study Institute, Urbino, Italy, 1975. Volume 1 of Pergamon Marine Series. E. B. Kraus (Editor). Pergamon Press, Oxford, New York, N.Y., Toronto, Ont., Sydney, N.S.W., 1977, 333 pp., 22 tables, 99 figures, US $19.00, £11.25, ISBN 0-08-020611-5. Ecological Modelling, 4(1):73, January 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380078900510.

Nihoul:1979:SME

[Nih79] J. C. J. Nihoul. Simulation modelling of environmental problems: François N. Frenkiel and David W. Goodall (Editors). John Wiley, 1978, xv + 112 pp., 10 figures, £4.95, ISBN 0-471-99580-0. Ecological Modelling, 6(4):335, May 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900498.

Nguyen:1979:MSA

[NW79] V. Van Nguyen and Eric F. Wood. On the morphology of summer algae dynamics in non-stratified lakes. Ecological Modelling, 6 (2):117–131, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380079900309.

Nyholm:1978:SMP

[Nyh78] Niels Nyholm. A simulation model for phytoplankton growth and nutrient cycling in eutrophic, shallow lakes. Ecological Modelling, 4 (2–3):279–310, February 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/030438007890011X.

Olufeagba:1975:BVA

[OFA75] B. J. Olufeagba, R. H. Flake, and N. E. Armstrong. A boundary value approach for estuarine water quality modelling with results for Jamaica Bay, New York. Ecological Modelling, 1(1): 3–30, May 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900034.

ONeill:1979:AEE

[OR79] R. V. O’Neill and B. Rust. Aggregation error in ecological models. Ecological Modelling, 7(2):91–105, July 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900012.

Parker:1975:EPE

[Par75] Richard A. Parker. Environmental periodicity and ecosystem stability. Ecological Modelling, 1(2):91–103, July 1975. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380075900253.

Parker:1978:NRC

[Par78a] R. A. Parker. Nutrient recycling in closed ecosystem models. Ecological Modelling, 4(1):67–70, January 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900492.

Parker:1978:SPN

[Par78b] Richard A. Parker. Spatial patterns in a nutrient-plankton model. Ecological Modelling, 4(4):361–370, May 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900285.

Perkins:1978:SSI

[Per78] Richard J. Perkins. A strategy for species introduction into evolving model ecosystems. Ecological Modelling, 5(1):1–15, June 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380078900145.

Peterman:1977:GEE

[Pet77] Randall M. Peterman. Graphical evaluation of environmental management options: Examples from a forest-insect pest system. Ecological Modelling, 3(2):133–148, May 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380077900035.

Penumalli:1976:LSS

[PFF76] B. R. Penumalli, R. H. Flake, and E. Gus Fruh. Large scale systems approach to estuarine water quality modelling with multiple constituents. Ecological Modelling, 2(2):101–115, June 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380076900284.

Phipps:1979:SWF

[Phi79] Richard L. Phipps. Simulation of wetlands forest vegetation dynamics. Ecological Modelling, 7(4):257–288, October 1979. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380079900383.

Pechan:1975:REM

[PLM75] Edward H. Pechan, Ralph A. Luken, and Hugh M. Maynard. Relating environmental modeling to policy issues and decisions. Ecological Modelling, 1(1):49–65, May 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380075900058.

Padgett:1979:SMP

[PP79] W. J. Padgett and A. S. Papadopoulos. Stochastic models for prediction of BOD and DO in streams. Ecological Modelling, 6(4): 289–303, May 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900425.

Rennie:1979:NDA

[Ren79] G. C. Rennie. New directions in the analysis of ecological systems, parts 1, 2: G. S. Innis (Editor). Simulation Councils Proceedings Series, Volume 5 Numbers 1, 2. Simulation Councils Inc., La Jolla, California, 1977, 248 pp., US $20 per part or $35 for both. Ecological Modelling, 7(2):163, July 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900061.

Reynolds:1977:SAW

[Rey77] Peter J. Reynolds. Systems approach to water management: A. K. Biswas (Editor). McGraw-Hill Publishing Company, New York, N.Y., 1976, 450 pp., $23.00. Ecological Modelling, 3(1):73–76, February 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900254.

Rowland:1979:MWC

[RM79] C. M. Rowland and A. R. McLellan. A model of a wasp colony population, Paravespula vulgaris (L.). Ecological Modelling, 7(2): 151–162, July 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438007990005X.

Scavia:1976:UMC

[SE76] Donald Scavia and Brian J. Eadie. The use of measurable coefficients in process formulations — zooplankton grazing. Ecological Modelling, 2(4):315–319, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380076900144.

Seip:1979:MMU

[Sei79] Knut Lehre Seip. A mathematical model for the uptake of heavy metals in benthic algae. Ecological Modelling, 6(3):183– 197, March 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900127.

Shahane:1976:IMW

[Sha76] Ashok N. Shahane. Interdisciplinary models of water systems. Ecological Modelling, 2(2):117–145, June 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380076900296.

Sjoberg:1977:PSI

[Sjö77] Stig Sjöberg. Are pelagic systems inherently unstable? A model study. Ecological Modelling, 3(1):17–37, February 1977. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304380077900229.

Skogerboe:1979:GSM

[Sko79] Gaylord V. Skogerboe. Grassland simulation model. Ecological studies, volume 26: George S. Innis (Editor), Springer-Verlag, Berlin–Heidelberg–New York, 1978, 298 pp., DM 49.60, U.S. $24.80, ISBN 3-540-90269-4. Ecological Modelling, 7(4):311–312, October 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900401.

Seip:1979:MMD

[SLM^{+}79] Knut Lehre Seip, Gulbrand Lunde, Sigurd Melsom, Even Mehlum, Angelika Melhuus, and Hans Martin Seip. A mathematical model for the distribution and abundance of benthic algae in a Norwegian fjord. Ecological Modelling, 6(2):133–166, February 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900310.

Smith:1977:EDC

[Smi77] Robert H. Smith. Ecological diversity: E. C. Pielou. John Wiley and Sons, New York, N.Y., 1975, 165 pp., £8.90, US $16.00, ISBN 0-471-68925-4. Ecological Modelling, 3(4):316–317, October 1977. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380077900138.

Scavia:1976:DSC

[SP76] Don Scavia and Richard A. Park. Documentation of selected constructs and parameter values in the aquatic model cleaner. Ecological Modelling, 2(1):33–58, March 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380076900053.

Straskraba:1979:NCM

[Str79] Milan Straškraba. Natural control mechanisms in models of aquatic ecosystems. Ecological Modelling, 6(4):305–321, May 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900437.

Sorrentino:1975:EIR

[SW75] John A. Sorrentino and Andrew B. Whinston. The economic implications of recycling exhaustible natural resources: the case for crude oil. Ecological Modelling, 1(3):219–233, September 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380075900198.

Swartzman:1979:SMM

[Swa79] Gordon L. Swartzman. Simulation modeling of material and energy flow through an ecosystem: Methods and documentation.

Ecological Modelling, 7(1):55–81, June 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900097.

Tiwari:1978:SSD

[THR^{+}78] Jawahar Lal Tiwari, John E. Hobbie, James P. Reed, Donald W. Stanley, and Michael C. Miller. Some stochastic differential equation models of an aquatic ecosystem. Ecological Modelling, 4(1): 3–27, January 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900443.

Ulanowicz:1978:EME

[UFHH78] Robert E. Ulanowicz, David A. Flemer, Donald R. Heinle, and Rogers T. Huff. The empirical modeling of an ecosystem. Ecological Modelling, 4(1):29–40, January 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900455.

VanBavel:1976:DSW

[VA76] C. H. M. Van Bavel and J. Ahmed. Dynamic simulation of water depletion in the root zone. Ecological Modelling, 2(3):189–212, September 1976. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380076900223.

Vincent:1977:CTM

[VLG77] Thomas L. Vincent, Cho Seng Lee, and B. S. Goh. Control targets for the management of biological systems. Ecological Modelling, 3 (4):285–300, October 1977. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304380077900096.

Vanderploeg:1979:CUS

[VS79] Henry A. Vanderploeg and Donald Scavia. Calculation and use of selectivity coefficients of feeding: Zooplankton grazing. Ecological Modelling, 7(2):135–149, July 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380079900048.

Wagenet:1979:ATH

[WGWJ79] R. J. Wagenet, W. J. Grenney, G. L. Wooldridge, and J. J. Jurinak. An atmospheric–terrestrial heavy metal transport model.

I. Model theory. Ecological Modelling, 6(3):253–272, March 1979. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380079900152.

Walters:1975:CSB

[WHP75] Carl J. Walters, Ray Hilborn, and Randall Peterman. Computer simulation of barren-ground caribou dynamics. Ecological Modelling, 1(4):303–315, December 1975. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380075900137.

Whitaker:1976:PMA

[WM76] Gene A. Whitaker and Richard H. McCuen. A proposed methodology for assessing the quality of wildlife habitat. Ecological Modelling, 2(4):251–272, December 1976. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380076900089.

Wheeler:1978:SML

[WRR78] G. L. Wheeler, G. L. Rolfe, and K. A. Reinbold. A simulation model for lead movement in a watershed. Ecological Modelling, 5 (1):67–76, June 1978. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380078900170.

Wang:1978:MMD

[WVW78] Lawrence K. Wang, David Vielkind, and Mu Hao Wang. Mathematical models of dissolved oxygen concentration in fresh water. Ecological Modelling, 5(2):115–123, July 1978. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380078900340.

Yezzi:1979:DMN

[YU79] D. J. Yezzi and A. P. Uzzo. Dynamic model of nutrification in Huntington Bay, New York. Ecological Modelling, 6(1):59– 75, January 1979. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380079900589.

Zucchetto:1975:EET

[Zuc75] James Zucchetto. Energy-economic theory and mathematical models for combining the systems of man and nature, case study: the urban region of Miami, Florida. Ecological Modelling, 1(4):241– 268, December 1975. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380075900101.